Today Jason is 5 years older than his brother tim. 2 years ago tim's age was 3/4 of Jason's age. What are their ages?
Answer by stanbon(75887)

Today Jason is 5 years older than his brother tim. 2 years ago tim's age was 3/4 of Jason's age. What are their ages?
-----
Equations:
j = t + 5
t-2 = (3/4)(j-2)
---------------------
Subtitute for "j" and solve for "t":
t-2 = (3/4)(t+5-2)
-----
t-2 = (3/4)(t+3)
-----
4t-8 = 3t+9
t = 17 years (tim's age now)
----
solve for "j"
Jason = j = 17+5 = 22 years
=================================
Cheers,
Stan H.
==========================